OpenStudy (anonymous):

1 + cos (4x) + cos(8x) + cos(12x)....+ (cos nx).

OpenStudy (vishweshshrimali5):

There is a direct formula for this: \[\large{\sum_{i = 0}^{n} \cos{(2ix)} = \csc{x}\sin{((n+1)x)}\cos{(nx)}}\]

OpenStudy (vishweshshrimali5):

In this question, you have to replace n by n/2.

OpenStudy (vishweshshrimali5):

@megprowder will this help you ?

OpenStudy (anonymous):

thx

OpenStudy (vishweshshrimali5):

Np

mathslover (mathslover):

@megprowder - If you think @vishweshshrimali5 helped you, then you can give him a medal. You can also fan him and write a testimonial in order to appreciate his work. Also, welcome to OpenStudy! :)